How many ml of 10.0 M NaOH do I need to add to 100 ml of 1.00 M NaOH to get a solution that is 5.00 in NaOH?
1
Expert's answer
2020-10-12T13:26:57-0400

Let x ml of 10.0 M NaOH have to added, with the 100 ml of 1.00 M NaOH solution.


So, V = x ml, S = 10.0 M

& V' = 100 ml , S' = 1.00 M


Total volume of solution = (100 +x)ml



Hence the resulting concentration of NaOH =[(VS+V'S')/total volume] M

 =[(10x+200) /(100+x)] M



Given resulting concentration of NaOH = 5.00 M 


So, 

 [(10x+200) /(100+x)] = 5

Or, 10x + 200 = 500 + 5x

Or, 5x = 300

Or, x = 60


Hence, 60 ml of 10.0 M NaOH need to add to get the solution of required strength.